# HPSS: Heuristic Prompting Strategy Search for LLM Evaluators

### v1

@@ 4.2 Experimental Results
Human Alignment
The main results are presented in Table 2 and 3. Firstly, with the same generation times, HPSS substantially improves the performance of LLM evaluators compared to the starting point (i.e., the prompting strategy from MT-Bench), resulting in an average 29.4% relative improvement in Spearman correlation across various tasks and outperforming other Non-LLM and human-designed LLM evaluators by a large margin. Even with only 5% of the generation times, HPSS remains superior to G-Eval and CloserLook. Additional experiments in Appendix J also demonstrate that the Qwen2.5-14B-Instruct evaluator with HPSS substantially surpasses the human-designed Qwen2.5-72B-Instruct evaluator. Secondly, HPSS achieves a significant performance gain over other automatic prompt optimization methods, showing its better adaptability to NLG evaluation tasks. Finally, HPSS brings consistent performance improvements to different evaluation models, highlighting its cross-model generalization capability. We also validate that integrating HPSS with inference-time methods such as self-consistency Wang et al. (2023b) can further improve the performance of LLM evaluators in Appendix K.
@@
Ablation Study
In Table 2 and 3, we provide an ablation study on key components of HPSS for the Qwen-2.5-14B-Instruct evaluator, including the mutation type exploitation, the additional exploration term, and the entire heuristic-function-guided mutation mechanism. The results validate that all components contribute to the final performance of HPSS. Notably, removing the entire mechanism leads to the most significant performance degradation, highlighting its effectiveness. Moreover, the performance gains from the heuristic search mechanism vary across different datasets, probably due to different search difficulties.
@@
4.3 Analysis
The generalizability of prompting strategies across datasets.
@@ 4.3 Analysis
We calculate the mean and standard deviations of the advantages for each factor at the end of HPSS across different datasets and evaluation aspects. Table 5 shows that the advantages exhibit considerable variance across tasks and evaluators, indicating the necessity of finding appropriate prompting strategies tailored to specific tasks and evaluators. However, there are some common characteristics: (1) A moderate scoring scale of 1-10 generally enhances evaluator performance, whereas a too coarse-grained scoring scale of 1-3 is less effective. (2) Under greedy decoding, directly generating scores and employing human-written evaluation criteria without AutoCoT, Metrics, and Reference generally improves evaluation performance. (3) Regarding the placement order of different components, positioning the task description at the beginning seems to yield better results, possibly because this arrangement is more logically coherent and helps evaluators concentrate more on sample information and evaluation criteria. These findings provide insights into future evaluation prompt design. We also provide a case study of HPSS in Appendix L.
@@ 5 Conclusion
In this work, we integrate 8 key factors for the prompt of LLM evaluators and propose HPSS, a heuristic search method to automatically optimize the prompting strategy for adjusting these factors. HPSS leverages the expected performance advantages of each value as the heuristic function to guide the search process. Extensive experiments on four NLG evaluation tasks demonstrate the superiority of HPSS, yielding consistent performance gains over both human-designed evaluation prompts and automatic prompt optimization methods. Additionally, we analyze the characteristics of these factors and provide new insights into future prompt design for LLM evaluators.
@@ 6 Limitations
1) HPSS requires iterative search on a validation dataset with human annotations, introducing additional annotation and inference overheads. However, we believe that it is not a severe problem for the following three reasons: Firstly, the overall cost is affordable in the vast majority of scenarios. Under the default setting, where 50% of the dataset is selected as the validation dataset, the search cost for a specific evaluation aspect of one dataset is approximately \$8 for GPT-4o-mini, and approximately 40 minutes using 4 H100 GPUs for Qwen2.5-14B-Instruct (See Appendix G). Secondly, HPSS remains effective in low-resource scenarios. As shown in Figure 3, even when annotated data is scarce (reducing the validation dataset size to 10% of the entire dataset, $\sim$ 100 samples per dataset) , HPSS can still significantly outperform manually designed prompt templates. Finally, HPSS is efficient during inference on the test dataset, requiring only a single greedy decoding pass for each test sample while still achieving better performance than much larger manually designed LLM evaluators that require 20 generations and self-consistency decoding (See Table 2, Table 3, and Appendix J). Nonetheless, introducing well-designed early stopping strategies could potentially reduce the search overheads of HPSS, which is considered as important future work.
@@
2) HPSS mainly focuses on the prompting strategy optimization of the input prompts for LLM evaluators. Despite the importance of input prompts, other modules such as decoding and interaction strategies may also benefit the performance of LLM evaluators. Regarding decoding strategies, our experiments demonstrate that incorporating self-consistency decoding can enhance the performance of HPSS (See Appendix K). Regarding multiple LLM evaluators’ interaction, in our preliminary experiments, we follow the settings of ChatEval Chan et al. (2024) to explore the impact of interaction strategies with multiple Qwen2.5-14B-Instruct evaluators, using the prompting strategies found by HPSS. We find that different interaction strategies have negligible impacts on the final evaluation performance, which is similar to the self-consistency method that directly averages the evaluation results of multiple individual LLM evaluators. Given that the search space increases significantly in multi-agent interaction scenarios and that it is difficult for LLM evaluators with different output formats to interact directly, we leave the optimization of interaction strategies among multiple LLM evaluators as important future work.
Appendix A List of Evaluation Prompt Templates
This section lists all prompt templates applied throughout this study, including the prompt templates utilized to generate the final rating (Table 9, 11, 13, 15) and the templates used to generate Reference, AutoCoT, and Metrics (Table 10, 12, 14, 16). For these prompt templates, we generally refer to the reference-free single answer pointwise grading prompt from MT-Bench Zheng et al. (2023). However, due to the contents of some components of this template being mixed, we make minor adjustments to the order of some sentences to ensure that the search for factor Order can be conducted. We use the prompt template of LLMBar Zeng et al. (2024) to generate Metrics.
Appendix B Preliminary Experiment on Factor Effects
Appendix C Details of In-Context Example Selection
Appendix D Details of Benchmarks
Appendix E Metric Calculation
E.1 Pointwise Grading
@@ E.2 Pairwise Comparison
Appendix F Implementation Details of Baselines
@@
Appendix G Implementation Details of HPSS
@@
Under the computational budget described in Section 4 (i.e., 71), the search cost of HPSS (i.e., the inference times of the LLM evaluator) for a specific evaluation aspect of one dataset is approximately 70 times the size of the validation dataset. When the size of the validation dataset is 50% of the entire dataset, for GPT-4o-mini, the cost of HPSS for a specific evaluation aspect of one dataset is approximately \$8. In total, the overall cost of HPSS across all the datasets and evaluation aspects is approximately \$140. For Qwen2.5-14B-Instruct, the runtime of HPSS for a specific evaluation aspect of one dataset is approximately 40 minutes using 4 H100 GPUs and the vllm Kwon et al. (2023) inference framework, while the overall runtime of HPSS across all the datasets and aspects is approximately 12 hours. The overall costs for other automatic prompt optimization methods for LLM evaluators are the same as HPSS, which is affordable in the vast majority of scenarios. We also validate that even with 1/5 of the above search cost (reducing the validation dataset size to 10% of the entire dataset), HPSS can still significantly outperform human-designed LLM evaluators in Figure 3.
Appendix H Details of Algorithm Implementation
Appendix I Experiments on Pairwise Comparison
I.1 Experimental Setup
@@ I.2 Main Results
Table 6: Accuracy of different prompting methods based on Qwen-2.5-14B-Instruct on pairwise comparison datasets including AUTO-J (Eval-P), LLMBar, and MT-Bench.
Model Method Summeval Topical-Chat SFHOT HANNA
Qwen-2.5-72B-Instruct MT-Bench 0.502 0.634 0.362 0.460
CloseLook + ICL 0.481 0.592 0.316 0.470
CloseLook + ICL 0.525 0.623 0.300 0.517
Qwen-2.5-14B-Instruct HPSS 0.560 0.697 0.431 0.519
Table 7: Average performance comparison of Qwen2.5-72B-Instruct evaluator and Qwen2.5-14B-Instruct evaluator on Summeval, Topical-Chat, SFHOT, and HANNA. †  indicates that the corresponding method employs 20 generations with self-consistency.
Method Summeval Topical-Chat SFHOT HANNA
CloseLook + ICL 0.532 0.660 0.402 0.555
HPSS 0.564 0.740 0.431 0.545
HPSS 0.569 0.763 0.453 0.578
@@
Appendix J Comparison with Stronger LLM Evaluator
As shown in Table 7, we compare the performance of the Qwen2.5-14B-Instruct evaluator using the prompting strategies obtained by HPSS with the Qwen2.5-72B-Instruct evaluator using human-designed prompting strategies. Notably, Qwen2.5-14B-Instruct with HPSS achieve significantly better evaluation performance than the human-designed Qwen2.5-72B-Instruct evaluator, even with only 5% of the generation times. These results demonstrate the efficiency of HPSS.
Appendix K Incorporating HPSS with Inference-Time Methods
@@
7:              Sample new candidate prompting strategy $\mathcal{T}^{new}$ in $T_{adj}$ based on Equation 3
@@
13:                  Select $\mathcal{T}^{max}$ with the highest overall advantage and $\mathcal{T}^{max}\notin T$ based on Equation 6
@@
21:                  Update advantage $A$ based on Equation 4 and 5
@@
Appendix L Case Study
In Table 17, we present some cases of prompting strategies explored by HPSS for specific tasks. We find that there is a significant performance gap between the best-performing and worst-performing prompting strategies, emphasizing the sensitivity of LLM evaluators to prompting strategy. Figure 7 and 8 illustrate the original evaluation prompts from MT-Bench alongside the optimized one found by HPSS, specifically focusing on the aspect Complexity in HANNA and the aspect Coherence in Topical-Chat. The prompting strategies of these prompts are shown in the "origin" and "best" rows in Table 17 for their respective datasets and aspects. We observe that some of the prompting strategies found by HPSS include values that are rarely considered in human-designed evaluation prompts. For instance, for the aspect Complexity in HANNA, the prompting strategy found by HPSS places the input content (IC) first and the task description (TD) last. For the aspect Coherence in Topical-Chat, evaluation criteria are not used in HPSS. These results demonstrate the limitations of manual prompt design and underscores the importance of automatic prompting strategy optimization.
Furthermore, to intuitively show the reason why HPSS achieves better evaluation performance than human-designed LLM evaluators, we provide two judgment generation cases for different LLM evaluators in Table 18 and 19. The corresponding evaluation prompts are shown in Figure 9 and 10. In the first case from the text summarization task presented in Table 18, HPSS provides a balanced assessment by analyzing both strengths and weaknesses of the summary, with emphasis on overall fluency. In contrast, MT-Bench and CloserLook + ICL overemphasize minor issues like typos while overlooking its advantages in overall fluency, and exhibit some hallucinations in judgment. In the second case from the story generation task presented in Table 19, HPSS conducts a systematic analysis of the story and correctly identifies both key plots that effectively convey emotions and overly idealized plots that weaken emotional delivery. In contrast, MT-Bench and CloserLook + ICL each overlook one of these two important points, resulting in less accurate evaluations. These observations demonstrate that HPSS improves the ability of LLM evaluators to conduct comprehensive evaluations and achieve a balanced assessment of strengths and weaknesses within the input sample.

### v2

@@ 4.2 Experimental Results
Human Alignment.
The main results are presented in Table 2 and 3. Firstly, with the same generation times, HPSS substantially improves the performance of LLM evaluators compared to the starting point (i.e., the prompting strategy from MT-Bench), resulting in an average 29.4% relative improvement in Spearman correlation across various tasks and outperforming other Non-LLM and human-designed LLM evaluators by a large margin. Even with only 5% of the generation times, as baseline methods require 20 generations with self-consistency while HPSS only requires a single greedy decoding pass, HPSS remains superior to G-Eval and CloserLook. Additional experiments in Appendix H also demonstrate that the smaller Qwen2.5-14B-Instruct and GPT-4o-mini evaluators with HPSS substantially surpass the larger human-designed Qwen2.5-72B-Instruct and GPT-4o evaluators, respectively. Secondly, HPSS achieves a significant performance gain over other automatic prompt optimization methods, showing its better adaptability to NLG evaluation tasks. Finally, HPSS brings consistent performance improvements to different evaluation models, highlighting its cross-model generalizability. We also validate that integrating HPSS with inference-time methods such as self-consistency Wang et al. (2023b) can further improve the performance of LLM evaluators in Appendix I.
@@
Ablation Study.
In Table 2 and 3, we provide an ablation study on key components of HPSS for the Qwen2.5-14B-Instruct evaluator, including the mutation type exploitation, the additional exploration term, and the entire heuristic-function-guided mutation mechanism. The results validate that all components contribute to the final performance of HPSS. Notably, removing the entire mechanism leads to the most significant performance degradation, highlighting its effectiveness. Moreover, the performance gains from the heuristic search mechanism vary across different datasets, probably due to different search difficulties.
@@
4.3 Analysis
The generalizability of prompting strategies across datasets.
The contribution of each factor.
Factor Average $\rho$ $\Delta$
HPSS 0.536 0.000
- w/o Scoring Scale 0.497 -0.039
- w/o In-Context Example 0.507 -0.029
- w/o Evaluation Criteria 0.508 -0.028
- w/o Reference 0.516 -0.020
- w/o Chain-of-Thought 0.500 -0.036
- w/o AutoCoT 0.518 -0.018
- w/o Metrics 0.527 -0.009
- w/o Order 0.492 -0.044
Table 5: Average performance of HPSS on Summeval, Topical-Chat, SFHOT, and HANNA when each factor is individually removed. Qwen2.5-14B-Instruct is employed as the evaluation model.
To investigate the contribution of each factor to the overall performance, we conduct an ablation study by independently removing each factor and rerunning HPSS. The average performance of Qwen2.5-14B-Instruct evaluator on four datasets is presented in Table 5. The results indicate that each factor contributes to the final performance of HPSS. However, the importance of these factors varies. Metrics (i.e., self-generated sample-specific metrics for good answers) appear to be the least important, whereas Order (i.e., the placement order of each component in the evaluation prompt) seems to be the most important. We leave the dedicated search space design to achieve a better trade-off between search overhead and performance as important future work.
@@ 4.3 Analysis
We calculate the mean and standard deviations of the advantages for each factor at the end of HPSS across different datasets and evaluation aspects. Table 6 shows that the advantages exhibit considerable variance across tasks and evaluators, indicating the necessity of finding appropriate prompting strategies tailored to specific tasks and evaluators. However, there are some common characteristics: (1) A moderate scoring scale of 1-10 generally enhances evaluator performance, whereas a too coarse-grained scoring scale of 1-3 is less effective. (2) Under greedy decoding, directly generating scores and employing human-written evaluation criteria without AutoCoT, Metrics, and Reference generally improves evaluation performance. (3) Regarding the placement order of different components, positioning the task description at the beginning seems to yield better results, possibly because this arrangement is more logically coherent and helps evaluators concentrate more on sample information and evaluation criteria. These findings provide insights into future evaluation prompt design.
@@
Analysis on reward hacking.
Reward hacking is a widespread problem in the optimization process Skalse et al. (2022). To analyze this problem in the HPSS framework, we follow AlignBench Liu et al. (2024a) and introduce a new metric, Pairwise Agreement, to provide an alternative perspective on evaluator performance during our optimization process. Specifically, for different model-generated texts to the same query (e.g., different generated summaries of the same article for SummEval), we consider all possible pairs and convert their human-judge and LLM-judge scores into pairwise comparison results (i.e., win, tie, or lose), respectively. Then, we measure the agreement between LLM evaluators and humans. On the validation dataset of Summeval and HANNA, the average performance of the Qwen2.5-14B-Instruct evaluator under different search steps of HPSS are shown in Figure 5. The results demonstrate that the Spearman correlation and pairwise agreement simultaneously improve during the optimization process and verify that the optimization process of HPSS does not introduce reward hacking but genuinely enhances the performance of LLM evaluators. We also provide a case study of HPSS in Appendix K.
Figure 5: Spearman Correlation and Pairwise Agreement of Qwen2.5-14B-Instruct across different HPSS search steps on the validation datasets of Summeval and HANNA. The first 21 steps are in Initialization, which may cause fluctuations in the metrics.
@@ 5 Conclusion
In this work, we integrate 8 key factors for the prompting strategy of LLM evaluators and propose HPSS, a heuristic search method to automatically optimize the prompting strategy for adjusting these factors. HPSS leverages the expected performance advantages of each value as the heuristic function to guide the search process. Extensive experiments on four NLG evaluation tasks demonstrate the superiority of HPSS, yielding consistent performance gains over both human-designed evaluation prompts and automatic prompt optimization methods. Additionally, we validate the generalizability of prompting strategies found by HPSS and analyze the characteristics of these factors. Our experimental results may provide new insights into future prompt design for LLM evaluators.
@@ 6 Limitations
1) HPSS requires iterative search on a validation dataset with human annotations, introducing additional annotation and inference overheads. However, we believe that it is not a severe problem for the following three reasons: Firstly, the overall cost is affordable in the vast majority of scenarios. Under the default setting, where 50% of the dataset is selected as the validation dataset, the search cost for a specific evaluation aspect of one dataset is approximately \$8 for GPT-4o-mini, and approximately 40 minutes using 4 H100 GPUs for Qwen2.5-14B-Instruct (See Appendix F.3). Secondly, HPSS remains effective in low-resource scenarios. As shown in Figure 3, even when annotated data is scarce (reducing the validation dataset size to 10% of the entire dataset, $\sim$ 100 samples per dataset) , HPSS can still significantly outperform manually designed prompt templates. Finally, HPSS is efficient during inference on the test dataset, requiring only a single greedy decoding pass for each test sample while still achieving better performance than human-designed LLM evaluators that require 20 generations with self-consistency decoding and much larger human-designed LLM evaluators (See Table 2, Table 3, and Appendix H). Nonetheless, introducing well-designed early stopping strategies could potentially reduce the search overheads of HPSS, which is considered as important future work.
@@
2) HPSS mainly focuses on the prompting strategy optimization of the input prompts for LLM evaluators. Despite the importance of input prompts, other modules such as decoding and interaction strategies may also benefit the performance of LLM evaluators. Regarding decoding strategies, our experiments demonstrate that incorporating self-consistency decoding can enhance the performance of HPSS (See Appendix I). Regarding multiple LLM evaluators’ interaction, in our preliminary experiments, we follow the settings of ChatEval Chan et al. (2024) to explore the impact of interaction strategies with multiple Qwen2.5-14B-Instruct evaluators, using the prompting strategies found by HPSS. We find that different interaction strategies have negligible impacts on the final evaluation performance, which is similar to the self-consistency method that directly averages the evaluation results of multiple individual LLM evaluators. Given that the search space increases significantly in multi-agent interaction scenarios and that it is difficult for LLM evaluators with different output formats to interact directly, we leave the optimization of interaction strategies among multiple LLM evaluators as important future work.
Acknowledgements
This work was supported by the National Science Foundation for Distinguished Young Scholars (with No. 62125604). We would also like to thank Zhipu AI for sponsoring the computational resources in this work.
Appendix A Metric Calculation
A.1 Pointwise Grading
@@ A.2 Pairwise Comparison
Appendix B List of Evaluation Prompt Templates
This section lists all prompt templates applied throughout this study, including the prompt templates utilized to generate the final rating (Table 11, 13, 15, 17) and the templates used to generate Reference, AutoCoT, and Metrics (Table 12, 14, 16, 18). For these prompt templates, we generally refer to the reference-free single answer pointwise grading prompt from MT-Bench Zheng et al. (2023). However, due to the contents of some components of this template being mixed, we make minor adjustments to the order of some sentences to ensure that the search for the factor Order can be conducted. We use the prompt template of LLMBar Zeng et al. (2024) to generate Metrics.
Appendix C Preliminary Experiment on Factor Effects
Appendix D Details of In-Context Example Selection
Appendix E Details of Benchmarks
Appendix F Implementation Details
F.1 Algorithm Implementation of HPSS
F.2 Implementation Details of Baselines
@@
F.3 Implementation Details of HPSS
@@
Under the computational budget described in Section 4 (i.e., 71), the search cost of HPSS (i.e., the inference times of the LLM evaluator) for a specific evaluation aspect of one dataset is approximately 70 times the size of the validation dataset. When the size of the validation dataset is 50% of the entire dataset, for GPT-4o-mini, the cost of HPSS for a specific evaluation aspect of one dataset is approximately \$8. In total, the overall cost of HPSS across all datasets is approximately \$140. For Qwen2.5-14B-Instruct, the runtime of HPSS for a specific evaluation aspect of one dataset is approximately 40 minutes using 4 H100 GPUs and the vllm Kwon et al. (2023) inference framework, while the overall runtime of HPSS across all datasets is approximately 12 hours. The overall costs for other automatic prompt optimization methods for LLM evaluators are the same as HPSS, which is affordable in the vast majority of scenarios. We also validate that even with 1/5 of the above search cost (reducing the validation dataset size to 10% of the entire dataset), HPSS can still significantly outperform human-designed LLM evaluators in Figure 3.
Appendix G Experiments on Pairwise Comparison
G.1 Experimental Setup
@@ G.2 Main Results
Table 7: Accuracy of different prompting methods based on Qwen2.5-14B-Instruct on pairwise comparison datasets including AUTO-J (Eval-P), LLMBar, and MT-Bench.
Model Method Summeval Topical-Chat SFHOT HANNA Average
Qwen2.5-72B-Instruct MT-Bench 0.502 0.634 0.362 0.460 0.490
CloseLook + ICL 0.481 0.592 0.316 0.470 0.465
Qwen2.5-14B-Instruct HPSS 0.560 0.697 0.366 0.519 0.536
GPT-4o MT-Bench 0.479 0.642 0.295 0.420 0.459
CloseLook + ICL 0.519 0.639 0.300 0.487 0.486
GPT-4o-mini HPSS 0.564 0.740 0.431 0.545 0.570
Table 8: Average performance comparison of HPSS and human-designed LLM evaluators on Summeval, Topical-Chat, SFHOT, and HANNA. †  indicates that the corresponding method employs 20 generations with self-consistency.
Method Summeval Topical-Chat SFHOT HANNA Average
CloseLook + ICL 0.532 0.660 0.402 0.555 0.537
HPSS 0.564 0.740 0.431 0.545 0.570
HPSS 0.569 0.763 0.453 0.578 0.591
@@
Appendix H Comparison with Stronger LLM Evaluator
As shown in Table 8, we compare the performance of smaller evaluators (i.e., Qwen2.5-14B-Instruct, GPT-4o-mini) using the prompting strategies obtained by HPSS with that of stronger evaluators (i.e., Qwen2.5-72B-Instruct, GPT-4o) using human-designed prompting strategies. Notably, the smaller evaluators with HPSS achieve significantly better evaluation performance than the human-designed stronger evaluators across different datasets. These results demonstrate the efficiency of HPSS.
Appendix I Incorporating HPSS with Inference-Time Methods
@@
7:              Sample new candidate prompting strategy $\mathcal{T}^{new}$ in $T_{adj}$ based on Equation 4 $\triangleright$ Exploration
@@
13:                  Select $\mathcal{T}^{max}$ with the highest overall advantage and $\mathcal{T}^{max}\notin T$ based on Equation 7 $\triangleright$ Exploitation
@@
21:                  Update advantage $A$ based on Equation 5 and 6 $\triangleright$ Advantage update
@@
Appendix J The Selection of Default Prompting Template
Prompt Template Summeval Topical-Chat SFHOT HANNA Average
MT-Bench Zheng et al. (2023) 0.432 0.562 0.346 0.464 0.451
G-Eval Liu et al. (2023) 0.436 0.579 0.315 0.466 0.449
Themis Hu et al. (2024b) 0.396 0.531 0.322 0.474 0.431
Table 10: Average performance of GPT-4o-mini evaluator under different prompting templates on Summeval, Topical-Chat, SFHOT, and HANNA.
We choose the prompt template of MT-Bench as the starting point because it has been widely used in previous works about LLM-as-a-Judge Doddapaneni et al. (2024); Zhang et al. (2025); Song et al. (2025). To verify the effectiveness of this prompt template, we compare the performance of the GPT-4o-mini evaluator under different prompt templates in Table 10. We find that the MT-Bench prompt template achieves comparable performance to the best-performed prompt template across various datasets, which validates the rationale for choosing it as the starting point.
@@
Appendix K Case Study
In Table 19, we present some cases of prompting strategies explored by HPSS for specific tasks. We find that there is a significant performance gap between the best-performing and worst-performing prompting strategies, emphasizing the sensitivity of LLM evaluators to the prompting strategy. Figure 8 and 9 illustrate the original evaluation prompts from MT-Bench alongside the optimized one found by HPSS, specifically focusing on the aspect Complexity in HANNA and the aspect Coherence in Topical-Chat. The prompting strategies of these prompts are shown in the "origin" and "best" rows in Table 19 for their respective datasets and aspects. We observe that some of the prompting strategies found by HPSS include values that are rarely considered in human-designed evaluation prompts. For instance, for the aspect Complexity in HANNA, the prompting strategy found by HPSS places the input content (IC) first and the task description (TD) last. For the aspect Coherence in Topical-Chat, evaluation criteria are not used in HPSS. These results demonstrate the limitations of manual prompt design and underscore the importance of automatic prompting strategy optimization.
Furthermore, to intuitively show the reason why HPSS achieves better evaluation performance than human-designed LLM evaluators, we provide two judgment generation cases for different LLM evaluators in Table 20 and 21. The corresponding evaluation prompts are shown in Figure 10 and 11. In the first case from the text summarization task presented in Table 20, HPSS provides a balanced assessment by analyzing both strengths and weaknesses of the summary, with emphasis on overall fluency. In contrast, MT-Bench and CloserLook + ICL overemphasize minor issues like typos while overlooking its advantages in overall fluency, and exhibit some hallucinations in judgment. In the second case from the story generation task presented in Table 21, HPSS conducts a systematic analysis of the story and correctly identifies both key plots that effectively convey emotions and overly idealized plots that weaken emotional delivery. In contrast, MT-Bench and CloserLook + ICL each overlook one of these two important points, resulting in less accurate evaluations. These observations demonstrate that HPSS improves the ability of LLM evaluators to conduct comprehensive evaluations and achieve a balanced assessment of strengths and weaknesses within the input sample.